### v2

# Direct Sampling from Conditional Distributions by Sequential Maximum Likelihood Estimations

Abstract. We can directly sample from the conditional distribution of any log-affine model. The algorithm is a Markov chain on a bounded integer lattice, and its transition probability is the ratio of the UMVUE (uniformly minimum variance unbiased estimator) of the expected counts to the total number of counts. The computation of the UMVUE accounts for most of the computational cost, which makes the implementation challenging. Here, we investigated an approximate algorithm that replaces the UMVUE with the MLE (maximum likelihood estimator). Although it is generally not exact, it is efficient and easy to implement; no prior study is required, such as about the connection matrices of the holonomic ideal in the original algorithm.
Key Words and Phrases. $A$-hypergeometric system, discrete exponential family, Gröbner bases, iterative proportional scaling, log-affine model, Markov chain Monte Carlo, Metropolis algorithm, rational maximum likelihood estimator, uniformly minimum variance unbiased estimator
2020 Mathematics Subject Classification Numbers. 62R01, 33C90, 33F99, 62H17, 65C05
1 Introduction
Consider a discrete sample with state space $[m]:=\{1,2,\ldots,m\}$ for a positive integer $m$. In this paper, we will discuss the discrete exponential families in statistics, which are also called log-affine models or toric models.
Definition 1.1.
Let $A=(a_{ij})$ $\in\mathbb{Z}^{d\times m}$ be a matrix of integers such that no row or column is zero vector and $(1,\ldots,1)\in{\rm rowspan}(A)$. Let $x\in\mathbb{R}^{m}_{>0}$. The log-affine model associated with the configuration matrix $A$ is the set of probability distributions
$\mathcal{M}_{A}:={\rm cl}\{p\in{\rm int}(\Delta_{m-1}):\log p\in\log x+{\rm rowspan}(A)\},$
where $\Delta_{m-1}$ is the standard $m$-dimensional simplex. If $x=1$, the model is called the log-linear model.
In this paper, we consider the closure, called the extended log-affine model by [11], Section 4.2.3, whose support may be at the boundary of the simplex. This extension ensures the existence of the maximum likelihood estimate (MLE), whether or not observed counts satisfy constraints to be positive.
The vector $p$ in Definition 1.1 may be parameterized as
$p_{j}=\frac{x_{j}}{Z(t,x)}\prod_{i\in[d]}t_{i}^{a_{ij}},\quad j\in[m]$ (1)
with unknown parameters $t\in\mathbb{R}_{>0}^{d}$ and the normalization constant $Z(t,x)$ such that $\sum_{j\in[m]}p_{j}=1$. A sample consisting of observations in which counts of the $j$-th state is $u_{j}$, $j\in[m]$ taken from the multinomial distribution specified by the probability mass function (1) follows the probability law
$\mathbf{P}(U=u,AU=b)=|u|!\prod_{j\in[m]}\frac{p_{j}^{u_{j}}}{u_{j}!}=\frac{|u|!}{\{Z(t,x)\}^{|u|}}\prod_{i\in[d]}t_{i}^{b_{i}}\prod_{j\in[m]}\frac{x_{j}^{u_{j}}}{u_{j}!},$ (2)
where the total number of counts is $|u|:=u_{1}+\cdots+u_{m}$. The conditional distribution given by complete minimal sufficient statistics $b$ in the affine semigroup $\mathbb{N}A:=\{Av:v\in\mathbb{N}^{m}\}$, where $\mathbb{N}$ is the set of non-negative integers, is
$\mathbf{P}(U=u|AU=b)=\frac{1}{Z_{A}(b;x)}\frac{x^{u}}{u!},\quad x^{u}:=\prod_{j\in[m]}x_{j}^{u_{j}},\quad u!:=\prod_{j\in[m]}u_{j}!.$ (3)
The support of this distribution is determined by the linear transformation
$\mathcal{F}_{b}(A):=\{v\in\mathbb{N}^{m}:Av=b\}.$ (4)
The set (4) is called the $b$-fiber associated to the configuration matrix $A$ with the sufficient statistics $b$. The normalization constant, or the partition function
$Z_{A}(b;x):=\sum_{v\in\mathcal{F}_{b}(A)}\frac{x^{v}}{v!}$ (5)
is called the $A$-hypergeometric polynomial, or the GKZ-hypergeometric polynomial, defined by Gel’fand, Kapranov, and Zelevinsky in the 1980’s. Note that the condition $(1,\ldots,1)\in{\rm rowspan}(A)$ in Definition 1.1 means that this polynomial is homogeneous. We adopt the convention $Z_{A}(b;x)=0$ if $b\notin\mathbb{N}A$. It is straightforward to see that a sample consisting of the Poisson random variables with means $\lambda_{j}$, $j\in[m]$ with $p_{j}=\lambda_{j}/|\lambda|$ also follows the conditional distribution (3).
Since the discrete exponential family is the standard family of discrete distributions, sampling from the conditional distribution (3) has various statistical demands. However, it also has been considered usually impossible (see, e.g. [20], Chapter 9). This motivates us to use the Metropolis algorithm, originally devised for sampling from canonical distributions in physics and one of the most common tools in the Markov chain Monte Carlo methods. In the Metropolis algorithm, the unique stationary distribution of an ergodic Markov chain is the distribution we need (see, e.g., [13], Chapter 3). We call the chain Metropolis chain. Diaconis and Sturmfels [6] proposed by using Gröbner bases of the toric ideal of a configuration matrix to construct a basis of moves in the Metropolis chain. They called such a basis, or a generating set for the toric ideal, a Markov Basis. The state space of the chain is the $b$-fiber (4), and the ratio of two monomials, which correspond to the current state and the proposed next state, is computed in each step of the chain.
However, the Metropolis algorithm has unavoidable drawbacks in principle. It cannot give independent samples, and we must wait for the convergence of the Metropolis chain to the stationary distribution, whose assessment is a difficult problem to answer. Therefore, it would be desirable, if possible, to obtain independent samples directly from the target distribution.
Importance sampling is an algorithm that draws samples that follow an easily generated distribution, and by weighting them based on the ratio to the target distribution, we regard them as samples drawn from the target distribution. Importance sampling can obtain independent samples and has no issue of convergence. Chen et al. [2] proposed a sequential importance sampling algorithm that draws contingency tables sequentially by each cell by introducing the notion that they called the Markov subbase. In practice, constructing the distribution to sample from is non-trivial and requires solving an integer or linear programming problem for each step.
Moreover, it remains true that importance sampling is not a direct sampling from the target distribution. The performance of an importance sampling is greatly affected by the choice of the weights and the distribution to sample from. Naturally, the optimal case is when weights are not necessary when sampled from the target distribution. Hence, a fundamental problem remained: is it possible to sample directly from the target distribution?
The answer is yes; we can sample from the conditional distribution (3) exactly using the holonomic ideal generated by a system of linear partial differential equations that the $A$-hypergeometric polynomial satisfies [14] (see Definition 2.4). The realizability is guaranteed by Gröbner bases of the holonomic ideal generated by the $A$-hypergeometric system. Throughout this paper, we call the sampling algorithm proposed in [14] the direct sampling algorithm. See the monograph [15] for an extensive discussion.
The direct sampling algorithm is a random walk or a Markov chain on a bounded integer lattice defined below. One sample path of the Markov chain is equivalent to one contingency table. Note that while the sequential draw of the contingency table is the same as in Chen et al. [2], as Markov chains the two are quite different and the units of the direct sampling are individual counts rather than cells. Figure 1 is the Hasse diagram of an example.
Figure 1: The Markov lattice for the matrix in Example 2.8 with the maximum $b={}^{t}(1,2,2,1)$. A sample path is shown by solid edges.
Definition 1.2 ([16]).
Consider a matrix $A=(a_{ij})\in\mathbb{Z}^{d\times m}$ satisfying the conditions in Definition 1.1 and a vector $b\in\mathbb{N}A$. The Markov lattice $\mathcal{L}_{A}(b)$ is the bounded integer lattice embedded in the affine semigroup $\mathbb{N}A$ equipped with the partial order
$\beta\in\mathbb{N}A~{}~{}\text{and}~{}~{}\beta-a_{j}\in\mathbb{N}A\quad\Rightarrow\quad\beta-a_{j}\prec\beta,$
and the maximum and the minimum are $b$ and 00, respectively. Here, $a_{j}$ denotes the $j$-th column vector of $A$.
As we will see in Section 2, Markov transitions occur between the adjacent elements of the Markov lattice with the Markov kernel represented by the transition probability
$P(\beta,\beta-a_{j};x):=\frac{Z_{A}(\beta-a_{j};x)}{Z_{A}(\beta;x)}\frac{x_{j}}{\deg(\beta)},\quad j\in[m]$ (6)
for the transition from $\beta$ to $\beta-a_{j}$, where $\deg(\beta)$ is the degree of $Z_{A}(\beta;x)$. The computation of the transition probabilities dominates the computational cost of the direct sampling algorithm. Evaluation of (6) following the definition of the $A$-hypergeometric polynomial (5) is practically impossible since the number of terms in the polynomial is the cardinality of the fiber (4) that increases rapidly with the total number of counts. In Section 4 of [16], the authors reported some results of numerical experiments of exact computations with rational number arithmetic on a computer algebra system. For two-way contingency tables of the independence model, the direct sampling algorithm demanded a time of about one-tenth time of the Metropolis algorithm when we count the effective sample size of samples taken from the Metropolis chain. On the other hand, for a model without independence, the direct sampling algorithm demanded a thousand times the time the Metropolis algorithm demanded. For the latter case, although the computation is exact, the computational burden would prohibit the practical use of the direct sampling algorithm.
We explain the reason behind this contrast in the computational costs. For the independence model, the transition probability (6) reduces to the product of the proportion of the row marginal count to the total number of counts and the proportion of the column marginal count to the total number of counts, a natural consequence of the independence. Therefore, (6) reduces to a rational number, and the computation is immediate. This argument was extended to decomposable graphical models in Section 3.2 of [16]. On the other hand, we have to compute (6) with some effort for a model without independence. More specifically, we need to compute the connection matrices of the Pfaffian system of the holonomic ideal (see Section 2.4). But unfortunately, the computation demands time growing exponentially with the number of rows or columns.
More generally, the main obstacle common to using the Markov basis and the direct sampling algorithm for general models is the computational cost of Gröbner bases. The Gröbner basis guarantees the realizability of these methods, however, Buchberger’s algorithm for computing Gröbner bases is impractical since it demands a huge computational cost (see, e.g., [4], Section 2.9). Therefore, in practice, we have to avoid using Gröbner bases. The challenge is to obtain Markov bases or connection matrices by some consideration using properties of each model, rather than resorting to Buchberger’s algorithm. A book-length treatment of this issue for the Markov bases is [1] and for the direct sampling algorithm is [15].
In this paper, we show that the transition probability (6) is the ratio of the uniformly minimum variance unbiased estimator (UMVUE) of the expected count to the total number of counts, where for a vector of observed counts $u\in\mathbb{N}^{m}$, we call $\mu:=\mathbf{E}U$ the vector of expected counts. In the final section of Chapter 1, entitled “Exact Methods,” Haberman [9] argued that the UMVUE was impractical for two main reasons: first, the UMVUE generally does not lie in the model manifold $\mathcal{M}_{A}$ (the toric variety defined by the matrix $A$); second, it is computationally intensive. In response to this, in the final paragraph of the chapter, she concluded, “Since exact methods are usually impractical, approximate methods must be tried. One such method, that of maximum likelihood, provides estimation and test procedures which involve simple computational methods and which have large sample properties. This method is the subject of the remainder of the monograph.” In this context, the significance of the UMVUE seems to have been overlooked, and the discussion of the MLE has dominated the literature.
In line with this trend, special MLEs have been attracting attention in recent years. Discrete statistical models for which the MLE admits an expression of a rational function of the observed counts are called rational, and they were characterized by [10] and [7]. Such an expression of the MLE is called the rational MLE. The MLE of the expected count of decomposable graphical models is given by a rational function of the observed counts (see, e.g., [11], Section 4.4), and it was shown that a log-linear model is rational if and only if it is a decomposable graphical model [8]. In this paper, we will see that the UMVUE and the MLE coincide if and only if a log-linear model is a decomposable graphical model.
We restate the observation above about the direct sampling algorithm as follows: if a model is log-linear and rational, then the computation of the transition probability (6) is immediate, and the direct sampling algorithm works efficiently and exactly. Otherwise, we may replace the UMVUE with the MLE and run the algorithm using existing methods to compute the MLE, such as the iterative proportional scaling algorithm, with the expense of bias of the MLE, that asymptotically disappears because of its consistency.
This paper is structured as follows. In Section 2, we introduce the details of the direct sampling algorithm proposed in [14] and show that the transition probability is the ratio of the UMVUE of the expected count to the total number of counts. Then, after summarizing the relevant properties of the MLE, we show that the UMVUE and the MLE coincide if and only if a log-linear model is a decomposable graphical model and the resulting MLE is rational. Lastly, we discuss the use of the MLE and reveal the key how we avoid computation of the connection matrices of the holonomic ideal. Section 3 is devoted to numerical experiments. Using several contingency tables as examples, we demonstrate how the direct sampling algorithm works. Section 4 summarizes the significance of this paper.
To the best of author’s knowledge, this is the first paper to demonstrate direct sampling from the conditional distribution of the discrete exponential family in a feasible time, although it is generally approximate in the sense that it is only asymptotically exact, not exact.
2 Algorithm
The direct sampling algorithm discussed in this paper is as follows.
Algorithm 1.
[14]
  • Input: A matrix $A\in\mathbb{Z}^{d\times m}$ satisfying the conditions in Definition 1.1, vectors $x\in\mathbb{R}_{>0}^{m}$, $b\in\mathbb{N}A$ with the total number of counts $n$.
  • Output: A vector of counts $u\in\mathbb{N}^{m}$ following the conditional distribution (3) given $b$.
  • Step 1: Set $t=1$, $\beta=b$, and $\nu=n$.
  • Step 2: Compute $P(\beta,\beta-a_{j};x)$ for each $j\in[m]$ in (6).
  • Step 3: Pick $j_{t}=j$ with probability $P(\beta,\beta-a_{j};x)$, $j\in[m]$.
  • Step 4: If $\nu=1$, go to Step 5. Else, set $t\leftarrow t+1$, $\beta\leftarrow\beta-a_{j_{t}}$, $\nu\leftarrow\nu-1$, and go to Step 2.
  • Step 5: Output $u_{j}:=\#\{t\in[n]:j_{t}=j\}$, $j\in[m]$.
In Input $b$ and $n$ should be consistent: since $(1,\ldots,1)\in{\rm rowspan}(A)$, we can always have the vector $c\in\mathbb{Q}^{d}$ with respect to $A$ such that $n=\sum_{i\in[d]}c_{i}b_{i}$.
Our concern is how to achieve Step 2, since the cost of this step dominates the computational cost of Algorithm 1. In the paper [14], the author obtained the connection matrices of the holonomic ideal for the case of $d=2$ and discussed computing $A$-hypergeometric polynomials to appear in advance. In the paper [16], the authors discussed computing $A$-hypergeometric polynomials needed along a sample path of the Markov chain for general cases. Both implementations give exact results, but the latter is general but not efficient.
Remark 2.1.
A merit of Algorithm 1 is that we do not have to construct the Markov lattice $\mathcal{L}_{A}(b)$. A Markov transition occurs only between adjacent elements because if otherwise, $Z_{A}(\beta-a_{j};x)=0$ since $\beta-a_{j}\not\in\mathbb{N}A$. For the same reason, a sample path never hits outside of $\mathcal{L}_{A}(b)$, or more precisely, the set $({\rm Cone}(A)\cap\mathbb{Z}^{m})\setminus\mathbb{N}A$, where ${\rm Cone}(A):=\{Av:v\in\mathbb{R}^{m}_{\geq 0}\}$ is the polyhedral cone generated by the column vectors of $A$.
2.1 The MLE
We introduce the MLE of the vector of expected counts. The probability law (2) leads to the log-likelihood
$l(\xi;b)=\sum_{i\in[d]}b_{i}\xi_{i}-\psi(\xi),\quad\psi(\xi):=|u|\log Z(e^{\xi},x)+{\rm const.},$
where $\xi_{i}=\log t_{i}$, $i\in[d]$ are the canonical parameters of the exponential family. The expected counts $\mu_{j}$, $j\in[m]$ satisfy
$\sum_{j\in[m]}a_{ij}\mu_{j}:=\sum_{j\in[m]}a_{ij}\mathbb{E}[U_{j}]=\mathbb{E}[B_{i}]=\frac{\partial\psi}{\partial\xi_{i}},\quad i\in[d],$
where the last equality holds by $\mathbb{E}[\partial l/\partial\xi_{i}]=0$. The MLEs of the canonical parameters $\hat{\xi}_{i}$, $i\in[d]$ constitute the unique solution of the estimating equations $b_{i}=\partial\psi/\partial\xi_{i}$. Therefore, we have
$\sum_{j\in[m]}a_{ij}\hat{\mu}_{j}=\left.\frac{\partial\psi}{\partial\xi_{i}}\right|_{\xi=\hat{\xi}}=b_{i},\quad i\in[d].$
More precisely, we have
Theorem 2.2 ([11], Theorem 4.8).
The MLE $\hat{\mu}$ of the vector of expected counts $\mu$ in the log-affine model $\mathcal{M}_{A}$ satisfies
$A\hat{\mu}=b$ (7)
for a sample of the vector of sufficient statistics $b$ is unique. This holds under multinomial as well as Poisson sampling.
In what follows, we sometimes denote the vector of the MLE $\hat{\mu}$ by $\hat{\mu}(b;x)$ to explicitly show the dependence on sufficient statistics $b$ and $x$. We apply this rule to the expected counts $\mu$ and the UMVUE $\tilde{\mu}$ discussed below.
Some components of the MLE become zero if sufficient statistics lie in the boundary of ${\rm Cone}(A)$. Related to Definition 1.2, we have the following proposition.
Proposition 2.3.
Consider a matrix $A=(a_{ij})\in\mathbb{Z}^{d\times m}$ satisfying the conditions in Definition 1.1 and a vector of sufficient statistics $b\in\mathbb{N}A$. If $b-a_{j}\not\in\mathbb{N}A$ for some $j$, then the component of the MLE $\hat{\mu}_{j}(b;x)$ vanishes.
Proof.
Since $b\in\mathbb{N}A$ and $b-a_{j}\not\in\mathbb{N}A$ for some $j$, if we write $b=\sum_{k\in[m]}v_{k}a_{k}$ for some $v\in\mathbb{N}^{m}$ (such $v$ is generally not unique), then $v_{j}=0$. This fact means that $b$ lies in a face of ${\rm Cone}(A)$, or
$b\in\{Av:v\in\mathbb{N}^{m},v_{j}=0\}\subset\{Av:v\in\mathbb{R}^{m}_{\geq 0},v_{j}=0\}.$
On the other hand, $\hat{\mu}$ satisfies (7), which implies $\hat{\mu}_{j}(b;x)=0$. ∎
In Section 2.4, we propose to rewrite Steps 2 and 3 as follows:
  • Step 2’: Compute the MLE of the expected count $\hat{\mu}_{j}(\beta;x)$ for each $j\in[m]$ of the log-affine model $\mathcal{M}_{A}$ with the sufficient statistics $\beta$;
  • Step 3’: Pick $j_{t}=j$ with probability $\hat{\mu}_{j}(\beta;x)/\nu$, $j\in[m]$,
and discuss the implementation.
2.2 The UMVUE and Markov transitions
We recall some basic properties of the $A$-hypergeometric polynomial (5).
Definition 2.4.
For a matrix $A=(a_{ij})\in\mathbb{Z}^{d\times m}$ and a vector $\beta\in\mathbb{C}^{d}$, the $A$-hypergeometric system is the following system of linear partial differential equations for an indeterminate function $f=f(x)$:
$(\sum_{j\in[m]}a_{ij}x_{j}\partial_{j}-\beta_{i})f=0,\quad i\in[d],\quad\partial_{j}:=\frac{\partial}{\partial x_{j}},\quad\text{and}$ (8)
$(\partial^{z^{+}}-\partial^{z^{-}})f=0,\quad z\in{\rm Ker}A\cap{\mathbb{Z}}^{m},\quad\partial^{z}:=\prod_{j\in[m]}\partial_{j}^{z_{j}},$ (9)
where $z^{+}_{j}=\max\{z_{j},0\}$ and $z^{-}_{j}=-\min\{z_{j},0\}$. The second group of operators in (9) generates the toric ideal of $A$.
We know that the $A$-hypergeometric polynomial with Definition 1.1 satisfies the homogeneity:
$(\sum_{j\in[m]}x_{j}\partial_{j}-\deg(\beta))Z_{A}(\beta;x)=0,$ (10)
and a simple computation shows the $A$-hypergeometric polynomial satisfies the contiguity relation:
$\partial_{j}Z_{A}(\beta;x)=Z_{A}(\beta-a_{j};x),\quad j\in[m].$ (11)
These properties give
Lemma 2.5.
The UMVUE $\tilde{\mu}$ of the vector of expected counts $\mu$ in the log-affine model $\mathcal{M}_{A}$ is given as
$\tilde{\mu}_{j}(b;x)=\frac{Z_{A}(b-a_{j};x)}{Z_{A}(b;x)}x_{j},\quad j\in[m],$ (12)
for $b\in\mathbb{N}A$, where $Z_{A}(b;x)$ is the $A$-hypergeometric polynomial defined by (5). Moreover, it is the unique unbiased estimator that is a function of $b$.
Proof.
By the conditional distribution (3), we observe that
$\mathbb{E}(U_{j}|AU=b)=\sum_{v\in\mathcal{F}_{b}(A)}v_{j}\frac{x^{v}}{v!}=\frac{\partial_{j}Z_{A}(b;x)}{Z_{A}(b;x)}x_{j}=\frac{Z_{A}(b-a_{j};x)}{Z_{A}(b;x)}x_{j},\quad j\in[m],$
where the last equality follows by the contiguity relation (11). This is the unbiased estimator of $\mu_{j}$ because $\mathbb{E}\{\mathbb{E}(U_{j}|AU=b)\}=\mathbb{E}[U_{j}]=\mu_{j}$. Since $b$ is the vector of complete sufficient statistics, by the Lehmann-Scheffé theorem (Theorem 2.1.11, [12]), it is the unique unbiased estimator that is a function of $b$. ∎
By the homogeneity (10), we have
$|u|=\sum_{j\in[m]}\tilde{\mu}_{j}(b;x)=\sum_{j\in[m]}\frac{x_{j}\partial_{j}Z_{A}(b;x)}{Z_{A}(b;x)}={\deg}(b).$
Hence, we conclude that the transition probability (6) is the ratio of the UMVUE of the expected counts to the total number of counts.
Before closing this subsection, we mention an observation about the UMVUE. Note that the following proposition does not mean that the UMVUE always coincides with the MLE since the UMVUE generally does not lie in the model manifold $\mathcal{M}_{A}$. If it lies, it should coincide with the MLE by Theorem 2.2. We will be back to this issue in Section 2.3.
Proposition 2.6.
The UMVUE $\tilde{\mu}$ of the vector of expected counts $\mu$ in the log-affine model $\mathcal{M}_{A}$ satisfies the estimating equation of the MLE (7), that is,
$A\tilde{\mu}=b$ (13)
for a sample of the vector of sufficient statistics $b$.
Proof.
Applying the annihilator (8) to the $A$-hypergeometric polynomial $Z_{A}(b;x)$ and using the contiguity relation (11), we have
$\sum_{j\in[m]}a_{ij}x_{j}Z_{A}(b-a_{j};x)=b_{i}Z_{A}(b;x),\quad i\in[d],$
which is equivalent to (13) due to the expression of the UMVUE (12). ∎
2.3 The UMVUE and the rational MLE
Step 2’ of Algorithm 1 is immediate if we know a closed form expression of the MLE in advance. In this respect, we have the following result based on the fact that the number of critical points of the likelihood over the complex space is one if and only if the map from the observed counts to the MLE admits a specific form called the Horn uniformization [10].
Theorem 2.7 ([7]).
Consider a sample of counts $u\in\mathbb{N}^{m}$ following the log-affine model $\mathcal{M}_{A}$. It has the rational MLE if and only if there is a vector $\lambda\in\mathbb{R}^{m}$ and a matrix $H=(h_{ij})\in\mathbb{Z}^{l\times m}$ such that $\mathcal{M}_{A}$ is the image of the rational map $\Phi:\mathbb{R}^{m}_{>0}\dasharrow\mathbb{R}^{m}_{>0}$:
$\Phi_{j}(u_{1},\ldots,u_{m})=\lambda_{j}\prod_{i=1}^{l}(\sum_{k=1}^{m}h_{ik}u_{k})^{h_{ij}},\quad j\in[m]$
with $\sum_{j\in[m]}\Phi_{j}(u)=1$.
Here, we present a small example in [10, 7] in our context.
Example 2.8 (Two-way contingency tables of the independence model).
We discuss the $2\times 2$ contingency table, but the same argument holds for general two-way contingency tables. The $2\times 2$ table is
$\begin{array}[]{cc|c}u_{11}&u_{12}&u_{1\cdot}\\ u_{21}&u_{22}&u_{2\cdot}\\ \hline\cr u_{\cdot 1}&u_{\cdot 2}&|u|\end{array},$
where $u_{i\cdot}=u_{i1}+u_{i2}$ and $u_{\cdot j}=u_{1j}+u_{2j}$, $i,j\in[2]$. The configuration matrix, the vectors of the observed counts, and the vector of the sufficient statistics are
$A\!=\!\left(\begin{array}[]{cccc}1&1&0&0\\ 0&0&1&1\\ 1&0&1&0\\ 0&1&0&1\end{array}\right),\quad u\!=\!\left(\begin{array}[]{c}u_{11}\\ u_{12}\\ u_{21}\\ u_{22}\end{array}\right),\quad{\rm and}\quad b=\left(\begin{array}[]{c}u_{1\cdot}\\ u_{2\cdot}\\ u_{\cdot 1}\\ u_{\cdot 2}\end{array}\right),$
respectively. The conditional distribution is
$\mathbf{P}(U=u|AU=b)=\frac{1}{Z_{A}(b;1)}\frac{1}{u_{11}!u_{12}!u_{21}!u_{22}!}.$
We have the rational MLE of the expected count $\hat{\mu}(b;1)=|u|\Phi(u)$ with
$\Phi(u)=\left(\frac{u_{1\cdot}u_{\cdot 1}}{|u|^{2}},\frac{u_{1\cdot}u_{\cdot 2}}{|u|^{2}},\frac{u_{2\cdot}u_{\cdot 1}}{|u|^{2}},\frac{u_{2\cdot}u_{\cdot 2}}{|u|^{2}}\right),$
and from this expression, we can read off
$\lambda=(4,4,4,4)\quad\text{and}\quad H=\begin{pmatrix}1&1&0&0\\ 0&0&1&1\\ -2&-2&-2&-2\\ 1&0&1&0\\ 0&1&0&1\end{pmatrix},$
where the rows are for $u_{1\cdot}$, $u_{2\cdot}$, $|u|$, $u_{\cdot 1}$, and $u_{\cdot 2}$, respectively. The Markov lattice $\mathcal{L}_{A}(b)$ with $b={}^{t}(1,2,2,1)$ is in Figure 1, and the transition probability of Step 3’ of Algorithm 1 becomes
$\frac{\hat{\mu}(b;1)}{|u|}=\frac{u_{i\cdot}}{|u|}\frac{u_{\cdot j}}{|u|},\quad i,j\in[2],$ (14)
where $a_{ij}$ denotes the column vector of $A$ cording the $(i,j)$-cell of the table.
In the remainder of this subsection we concentrate on the log-lineal model, so we set $x=1$.
Geiger et al. [8] showed that a log-linear model has the rational MLE if and only if it is a decomposable graphical model (see [11] for the definition of the decomposable graphical model). Using this fact, we have the following theorem.
The notation for graphical models can be complex; we follow that of [11] here. The correspondence between this notation and the notation of the $A$-hypergeometric polynomial is explained in detail in Section 3.2 of [16], so we will not repeat it here. The two-way contingency table of the independence model discussed above is an example of a decomposable graphical model: the undirected graph consists of two vertices corresponding to the row and column of the contingency table and no edge. The two vertices are the cliques, and the separator is empty.
Theorem 2.9.
The UMVUE and the MLE of a log-linear model coincides if and only if the model is a decomposable graphical model. Moreover, the MLE is rational.
Proof.
(if) Lemma 4.21 of [11] gives the normalization constant of a decomposable graphical model. We can represent it as the $A$-hypergeometric polynomial, and we have
$Z_{A}(b;1)=\frac{\prod_{S\in\mathcal{S}}\{\prod_{j_{S}}u(j_{S})!\}^{\nu(S)}}{\prod_{C\in\mathcal{C}}\prod_{j_{C}}u(j_{C})!},$
where $\mathcal{S}$ and $\mathcal{C}$ are the sets of separators and cliques of the undirected graph $G$, respectively, and $\nu(S)$ is the number of appearances of separator $S$ in a perfect sequence of the cliques. The index $j$ specifies a state of the vertex set of $G$. The indices $j_{S}$ and $j_{C}$ are indices obtained by summing up indices of $j$ other than $S$ and $C$, respectively, and $u(j_{S})$ and $u(j_{C})$ are the number of counts of states consistent with $j_{S}$ and $j_{C}$, respectively. Lemma 2.5 shows that the UMVUE is
$\tilde{\mu}_{j}(b;1)$ $=\frac{Z_{A}(b-a_{j};1)}{Z_{A}(b;1)}$
$=\frac{\prod_{S\in\mathcal{S}}\prod_{i_{S}\neq j_{S}}\{u(i_{S})!\}^{\nu(S)}\{(u(j_{S})-1)!\}^{\nu(S)}}{\prod_{C\in\mathcal{C}}\prod_{i_{C}\neq j_{C}}u(i_{C})!(u(j_{C})-1)!}\frac{\prod_{C\in\mathcal{C}}\prod_{i_{C}}u(i_{C})!}{\prod_{S\in\mathcal{S}}\prod_{i_{S}}\{u(i_{S})!\}^{\nu(S)}}$
$=\frac{\prod_{C\in\mathcal{C}}u(j_{C})}{\prod_{S\in\mathcal{S}}u(j_{S})^{\nu(S)}}.$
The right-hand side is the expression of the MLE $\hat{\mu}_{j}(b;1)$ given in Proposition 4.18 of [11] and the rational function of observed counts.
(only if) Suppose we have a nondecomposable graphical model. Then, Theorem 4.4 of [8] implies that we cannot always find a rational solution to the maximum likelihood estimation. On the other hand, the expression of the UMVUE in Lemma 2.5 implies that the UMVUE is always a rational number, because of the definition of the $A$-hypergeometric polynomial (5). Therefore, the MLE and the UMVUE never coincide. ∎
With Proposition 2.6, we immediately have
Corollary 2.10.
The UMVUE of a log-linear model lies in the model manifold $\mathcal{M}_{A}$ if and only if the model is a decomposable graphical model.
2.4 The approximate algorithm without using connection matrices
In this subsection, we propose to use Algorithm 1 with Steps 2 and 3 replaced by Steps 2’ and 3’, respectively. We call the resulting algorithm as the approximate direct sampling algorithm. From Lemma 2.4, when the MLE is not the UMVUE, i.e., it is not the rational MLE, then there exists a bias that disappears asymptotically. Thus the resulting algorithm is approximate in the sense that it is only asymptotically exact, not exact.
Before proceeding, we summarize what we know about Algorithm 1 according to Theorem 2.9 as a proposition.
Proposition 2.11.
For a log-linear model, the approximate algorithm obtained from Algorithm 1 by replacing Steps 2 and 3 with Steps 2’ and 3’ coincides with the original Algorithm 1 and thus becomes exact if and only if the model is a decomposable graphical model.
Theorem 2.7 means that the time computational complexity of Step 2 of Algorithm 1 is $O(lm)$ if we use the rational MLE.
Remark 2.12.
The fact that Step 2 of Algorithm 1 is immediate by the rational MLE for decomposable graphical models is not surprising, given that we can sample directly by combining urn schemes (see, e.g., [20], Section 9.6). However, an implementation of Algorithm 1 may be easier.
Combining Remark 2.1 and Proposition 2.3, we arrive at the following consequence: if sufficient statistics lie in the boundary of the polyhedral cone generated by the column vectors of $A$, the MLE coincides with the UMVUE and vanishes exactly without bias. In practical terms, we obtain
Proposition 2.13.
The approximate algorithm obtained from Algorithm 1 by replacing Steps 2 and 3 with Steps 2’ and 3’ is a Markov chain on the Markov lattice defined in Definition 1.2.
For any log-affine model, we can always evaluate the MLE numerically solving the estimating equation (7) by using the iterative proportional scaling (IPS), specifically called the generalized IPS by Darroch and Ratcliff [5]. We rewrite the procedure in our notation. Consider a sample of counts $u\in\mathbb{N}^{m}$ following the log-affine model $\mathcal{M}_{A}$ satisfying $Au=\beta$, where $\beta\in\mathbb{N}A$. For simplicity, we assume $A\in\mathbb{N}^{d\times m}$, $\beta_{i}>0,i\in[d]$, and all columns of $A$ have the same sum denoted by $s$, but a general form is available in [5]. That is
$p_{j}^{(0)}=\frac{x_{j}}{\sum_{k\in[m]}x_{k}},\quad j\in[m],\quad\text{and}$
$p_{j}^{(t+1)}=p_{j}^{(t)}\left\{\prod_{i\in[d]}\left(\frac{\beta_{i}/|u|}{\sum_{k\in[m]}a_{ik}p_{k}^{(t)}}\right)^{a_{ij}}\right\}^{1/s},\quad j\in[m].$ (15)
Theorem 1 of [5] guarantees the convergence: $\lim_{t\to\infty}|u|p^{(t)}=\hat{\mu}(\beta;x)$.
The time computational complexity of Step 2’ of Algorithm 1 is $O(dm\tau)$ if we use the generalized IPS, where $\tau$ is the minimum number of iterations to achieve convergence with predetermined precision. In the numerical experiments reported in Section 3, we used the iteration (15) with a convergence criterion to stop the iteration at the time
$\tau:=\min\left\{t\geq 1:\sum_{i=1}^{d}||u|\sum_{k\in[m]}a_{ik}p_{k}^{(t)}-\beta_{i}|<\epsilon d\right\}$ (16)
for some small $\epsilon>0$, which means the error per a sufficient statistic is less than $\epsilon$.
Remark 2.14.
There are various types of IPS algorithms for maximum likelihood estimation, and we may use any of them. The classical IPS known as the Deming–Stephan algorithm for log-linear models has the iteration with setting $s=1$ in (15). Coons et al. [3] considered a subfamily of the log-linear model and obtained a sufficient condition under which the classical IPS produces the rational MLE in one cycle of iteration.
We summarize the time computational complexity of the algorithms proposed in this paper. We only consider the time complexity because the space complexity, i.e., the memory requirement of the implementation, is small enough that they are of no concern.
Proposition 2.15.
The time computational complexity of Algorithm 1 is $O(lm|u|)$ if we use the rational MLE for Step 2, where $l$ is in Theorem 2.7. For the approximate algorithm obtained from Algorithm 1 by replacing Steps 2 and 3 with Steps 2’ and 3’, the complexity is $O(dm\tau|u|)$ if we use the generalized IPS for Step 2’. Here, (16) defines $\tau$.
In the approximate algorithm, we avoided the need to compute the connection matrices of the holonomic ideal generated by the $A$-hypergeometric system. Therefore, we can avoid the computation of Gröbner bases or investigation of specific properties of each model for any model. We explain how we achieved this before closing this section.
For the holonomic ideal, we have a Pfaffian system of equations:
$x_{j}\partial_{j}q(\beta;x)=F_{j}(\beta;x)q(\beta;x),\quad j\in[m]$ (17)
for some connection matrices $F_{j}(\beta;x)$, $j\in[m]$ and the vector
$q(\beta;x)={}^{t}(Z_{A}(\beta;x),\partial^{u(1)}Z_{A}(\beta;x),\ldots,\partial^{u(r-1)}Z_{A}(\beta;x)),$
where $\{1,\partial^{u(1)},\ldots,\partial^{u(r-1)}\}$ is the set of the standard monomials of the holonomic ideal of rank $r$, where $\partial^{u(\cdot)}$ is a monomial in $\partial_{j}$, $j\in[m]$. The standard monomials provide the basis of the solutions of the $A$-hypergeometric system. The system of equations (17) can be obtained via a normal form with respect to a Gröbner basis of the holonomic ideal (see, e.g., [18], Theorem 1.4.22). Then, with the contiguity relation (11), we obtain a recurrence relation for the vectors $q(\beta;x)$:
$x_{j}q(\beta-a_{j};x)=F_{j}(\beta;x)q(\beta;x),\quad j\in[m],$ (18)
which was originally used to compute the $A$-hypergeometric polynomial $Z_{A}(\beta;x)$ [17] and used in a previous implementation of the direct sampling algorithm ([16], Algorithm 2.2). Nevertheless, Lemma 2.5 is equivalent to a relation
$x_{j}Z_{A}(\beta-a_{j};x)=\tilde{\mu}_{j}(\beta;x)Z_{A}(\beta;x),\quad j\in[m].$
Comparing this with (18), we see that by appropriate basis transformation we can make the first row of the connection matrix to be $(\tilde{\mu}_{j}(\beta;x),0,\ldots,0)$. All we need is an evaluation of the UMVUE $\tilde{\mu}$ exactly or approximately, since the recurrence relation for the first element of the transformed vector decouples from the others. Here, the MLE asymptotically coincides with the UMVUE and can be used as an approximation of the UMVUE, but we can compute the MLE without using the recurrence relation.
3 Numerical experiments
We applied the direct sampling algorithm and the approximate algorithm introduced in Section 2.4 to three log-affine models. The first two models are two-way contingency tables. Since the independence model introduced in Example 2.8 is a decomposable graphical model, the UMVUE coincides with the rational MLE by Theorem 2.9, and the original Algorithm 1 works efficiently. The other is a non-independence model, which is not log-linear. In Section 4.2 of [16], the authors applied a previous implementation of the direct sampling algorithm (Algorithm 2.2 of [16]) to the non-independence model. It is exact, but they found it was inefficient due to the computational burden. The last model is the no-three-way interaction model for the three-way contingency table. It is log-linear, but not a decomposable graphical model, so the UMVUE and the MLE never coincide. In Section 4.3 of [16], the authors discussed the model but did not experiment since it involves the recurrence relation (18) for a vector of 81 dimensions, and its computation was prohibitive. Because of the above reasons, we applied the approximate direct sampling algorithm to the last two models.
We implemented the direct sampling and the Metropolis algorithms on a computer and compared their performance because they are equally easy to implement, and the Metropolis algorithm is the standard in practitioners.
To evaluate the performance of the Metropolis algorithm, it is first necessary to assess the convergence of the Metropolis chain to the stationary distribution. It is a nontrivial task with various arguments (see, e.g., [19], Section 12.2). In this study, taking advantage of the situation where we have another feasible sampling algorithm, we assessed the convergence of the Metropolis chain using the direct sampling algorithm.
Since the number of possible contingency tables is enormous, we employed the distribution of the chi-square values of the contingency table rather than directly assessing the distribution of contingency tables. We estimated the distribution of chi-square values under the conditional distribution (3) of contingency tables, that is,
$p(z):=\mathbf{P}(\chi^{2}(U)=z|AU=b),\quad z\geq 0,$ (19)
by the empirical distribution $\hat{p}_{\rm M}$ of contingency tables taken from the Metropolis chain and the empirical distribution $\hat{p}_{\rm D}$ of contingency tables taken by the direct sampling algorithm. We may regard these estimates as the Monte-Carlo integrations of the indicator function $1\{\chi^{2}(\cdot)=z\}$. Then, we computed the total variation distance between them:
$\|\hat{p}_{\rm M}-\hat{p}_{\rm D}\|_{\rm TV}=\frac{1}{2}\sum_{z}|\hat{p}_{\rm M}(z)-\hat{p}_{\rm D}(z)|^{2}.$ (20)
Here, we generated many contingency tables for $\hat{p}_{\rm D}$ in advance so that the empirical distribution of the direct sampling algorithm gives the best possible approximation of the truth (19). If we use the (exact) direct sampling algorithm, (20) approximates the total variation distance between the distribution of chi-square values taken from the Metropolis chain and that of the stationary distribution. If otherwise, we can still assess the convergence of the Metropolis chain because the total variation distance should converge to a non-zero value.
Contingency tables taken from the Metropolis chain are not independent. To account for the autocorrelation among contingency tables, we adopted the concept of the effective sample size (ESS). The ESS of $N$ consecutive steps in a Metropolis chain is $N/(1+2\sum_{t\geq 1}\rho_{t})$, where $\rho_{t}$ is the autocorrelation of chi-square values at lag $t$. We used sample autocorrelation and censored the sum where the sample autocorrelation was less than 0.01.
We conducted all computations in the R statistical computing environment on floating-point arithmetic. The following timing results were taken on a single core of the CPU (Intel Core i5-4308U CPU, 2.8GHz) of a machine with 8GB memory.
3.1 Two-way contingency tables
We investigated the performance of the direct sampling algorithm by exemplifying the following two-way contingency tables with fixed marginal sums for $s=1,2,5$ and $10$:
$\begin{array}[]{ccccc|c}u_{11}&u_{12}&u_{13}&u_{14}&u_{15}&5s\\ u_{21}&u_{22}&u_{23}&u_{24}&u_{25}&5s\\ u_{31}&u_{32}&u_{33}&u_{34}&u_{35}&5s\\ u_{41}&u_{42}&u_{43}&u_{44}&u_{45}&5s\\ \hline\cr 4s&4s&4s&4s&4s&20s\end{array},$
where the vector
$u={}^{t}(u_{11},\ldots,u_{15},u_{21},\ldots,u_{25},\ldots,u_{41},\ldots,u_{45}),$
is random, the vector of sufficient statistics is
$b={}^{t}(u_{1\cdot},\ldots,u_{4\cdot},u_{\cdot 1},\ldots,u_{\cdot 5})={}^{t}(5s,\ldots,5s,4s,\ldots,4s),$
and the configuration matrix is
$A=\left(\begin{array}[]{c}E_{4}\otimes 1_{5}\\ 1_{4}\otimes E_{5}\end{array}\right),\quad(e_{r})_{ij}:=\delta_{i,j},~{}i,j\in[r],\quad 1_{r}:=(\overbrace{1,\ldots,1}^{r}).$
These tables are unrealistically symmetric, but advantages of this symmetry for the following examinations are that we know the expected counts of each cell for the independent model is uniform, and we expect that the Metropolis chain starting from this tables will mix rapidly. The rational MLE of the expected count is (14) with $i\in[4]$ and $j\in[5]$. For the non-independence model, we set the odds ratios ($x$ in Definition 1.1) as
$\left(\begin{array}[]{ccccc}3&2&1&1&1\\ 2&2&1&1&1\\ 1&1&1&1&1\\ 1&1&1&1&1\\ \end{array}\right).$ (21)
For the generalized iterative proportional scaling (IPS), we adopt the convergence criterion (16) with $\epsilon=0.1$. All computations involved converged within 20 steps.
First, we generated one million contingency tables using the direct sampling algorithm (exact and approximate for the independence and non-independence models, respectively). We computed the empirical distribution $\hat{p}_{\rm D}$ of the chi-square values, which should be a good approximation to the stationary distribution for the independence model.
Then, we assessed the convergence of the Metropolis chain to the stationary distribution. For the Metropolis chain of four pairs of the burn-in and length: $(0,10^{3})$, $(10^{3},10^{4})$, $(10^{4},10^{4})$, and $(10^{4},10^{5})$, respectively, we computed the empirical distribution $\hat{p}_{\rm M}$ of the chi-square values and calculated the total variation distance (20). Here, we set the expected value of each cell in the chi-square to $s$ for both the independence and non-independence models with the odds ratios (21). We show the averages of the results obtained by repeating this procedure 10 times in Table 1. The pair $(0,10^{3})$ gave the poorest results because only a few contingency tables appeared in a Metropolis chain, and they cannot approximate the stationary distribution. The pairs $(10^{3},10^{4})$ and $(10^{4},10^{4})$ were equally good, which implies that the Metropolis chain converged after 10,000 steps. For the pair $(10^{4},10^{5})$, the approximation is even better because the empirical distribution approximates the stationary distribution better by many contingency tables because of the Glivenko–Cantelli theorem. These trends were similar for non-independence models. Another important observation is that for approximate sampling of the non-independence models, the total variation distance did not decrease as $s$ increases. While one would expect the total variation distance to decrease as the bias of the MLE decreases asymptotically, the reverse trend suggests that the Monte Carlo error is relatively large compared to the bias and that the impact of the bias is small in typical settings. Indeed, the total variation distances showed a similar trend for the direct sampling from the independence model.
Table 1: The total variation distances between the empirical distributions of chi-square values generated by the Metropolis algorithm and that by the direct sampling algorithm (exact and approximate for the independence and non-independence models, respectively).
(burn-in, length)
Model $s$ $(0,10^{3})$ $(10^{3},10^{4})$ $(10^{4},10^{4})$ $(10^{4},10^{5})$
Independence 1 0.133 0.039 0.035 0.012
2 0.137 0.041 0.043 0.015
5 0.192 0.055 0.062 0.019
10 0.236 0.072 0.073 0.025
Non-independence 1 0.107 0.036 0.031 0.017
2 0.141 0.047 0.046 0.025
5 0.190 0.063 0.065 0.029
10 0.273 0.089 0.083 0.037
Given the above results, we drew 10,000 tables after the burn-in of 10,000 steps for the Metropolis chain. We computed the ESS of the 10,000 tables drawn by the Metropolis algorithm and drew the same number of tables using the direct sampling algorithm. Table 2 shows the averages of the results obtained by repeating this procedure 100 times. It shows the times it took to draw 10,000 tables using the Metropolis algorithm (including the burn-in), the time it took to draw the tables of the number of the ESS using the direct sampling algorithm, and one table using the direct sampling algorithm. For the independence model, the direct sampling algorithm was uniformly efficient than the Metropolis algorithm. For the non-independence model, although the direct sampling algorithm was inefficient than the Metropolis algorithm, even when the total count is 200 ($s=10$), which is the most computationally expensive for the direct sampling algorithm, the direct sampling algorithm took only about six times as long as the Metropolis algorithm.
Table 2: The times to draw 10,000 tables by the Metropolis algorithm, the effective sample sizes (ESS) of the 10,000 tables, the times to draw the same number of tables by the direct sampling algorithm (exact and approximate for the independence and non-independence models, respectively), and the times per single table. Times are in seconds.
Model $s$ ESS Metropolis Direct 1 table
Independence 1 327 0.604 0.193 0.0006
2 265 0.663 0.308 0.0012
5 155 0.750 0.448 0.0029
10 103 0.839 0.594 0.0058
Non-independence 1 327 0.650 1.543 0.005
2 257 0.726 2.876 0.011
5 145 0.846 4.683 0.032
10 81 0.945 5.711 0.071
3.2 No-three-way interaction model of three-way table
We investigated the performance of the approximate direct sampling algorithm by exemplifying the following no-three-way interaction model of three-way contingency tables with fixed marginal sums for $s=1,2,5$ and $10$:
$\begin{array}[]{ccc|c}u_{111}&u_{112}&u_{113}&3s\\ u_{121}&u_{122}&u_{123}&3s\\ u_{131}&u_{132}&u_{133}&3s\\ \hline\cr 3s&3s&3s&9s\end{array}\quad\begin{array}[]{ccc|c}u_{211}&u_{212}&u_{213}&3s\\ u_{221}&u_{222}&u_{223}&3s\\ u_{231}&u_{232}&u_{233}&3s\\ \hline\cr 3s&3s&3s&9s\end{array}\quad\begin{array}[]{ccc|c}2s&2s&2s&6s\\ 2s&2s&2s&6s\\ 2s&2s&2s&6s\\ \hline\cr 6s&6s&6s&18s\end{array},$
where the vector
$u={}^{t}($ $u_{111},u_{112},u_{113},u_{121},u_{122},u_{123},u_{131},u_{132},u_{133},$
$u_{211},u_{212},u_{213},u_{221},u_{222},u_{223},u_{231},u_{232},u_{233}).$
is random, the vector of the sufficient statistics is
$b={}^{t}($ $u_{11\cdot},u_{12\cdot},u_{13\cdot},u_{21\cdot},u_{22\cdot},u_{23\cdot},u_{1\cdot 1},u_{1\cdot 2},u_{1\cdot 3},u_{2\cdot 1},u_{2\cdot 2},u_{2\cdot 3},$
$u_{\cdot 11},u_{\cdot 12},u_{\cdot 13},u_{\cdot 21},u_{\cdot 22},u_{\cdot 23},u_{\cdot 31},u_{\cdot 32},u_{\cdot 33})$
=($={}^{t}($ $3s,\ldots,3s,2s,\ldots,2s),$
and the configuration matrix is
$A=\left(\begin{array}[]{c}E_{6}\otimes 1_{3}\\ E_{2}\otimes 1_{3}\otimes E_{3}\\ 1_{2}\otimes E_{9}\end{array}\right).$
A Markov basis consists of [6]
$\begin{bmatrix}1i_{2}i_{3}\\ 1j_{2}j_{3}\\ 2i_{2}j_{3}\\ 2j_{2}i_{3}\end{bmatrix}-\begin{bmatrix}1i_{2}j_{3}\\ 1j_{2}i_{3}\\ 2i_{2}i_{3}\\ 2j_{2}j_{3}\end{bmatrix},\quad i_{2}\neq j_{2},~{}i_{3}\neq j_{3}\in[3],$
which represents nine vectors with the components of $u$ in the first bracket are $+1$, those in the seconds brackets are $-1$, and the other components are 0, and six vectors with 12 non-zero elements including
$\begin{bmatrix}11i_{1}\\ 12i_{2}\\ 13i_{3}\\ 21j_{1}\\ 22j_{2}\\ 23j_{3}\end{bmatrix}-\begin{bmatrix}11j_{1}\\ 12j_{2}\\ 13j_{3}\\ 21i_{1}\\ 22i_{2}\\ 23i_{3}\end{bmatrix},$
where $(i_{1},i_{2},i_{3})$ is a permutation of $(1,2,3)$ and determines $(j_{1},j_{2},j_{3})$. For example, if $(i_{1},i_{2},i_{3})=(1,3,2)$, then $(j_{1},j_{2},j_{3})=(2,1,3)$.
First, we assessed the convergence of the generalized IPS. Unlike the case of the two-way contingency tables in Section 3.1, setting $\epsilon$ in equation (16) needed care. Proposition 2.13 guarantees that a sample path is on the Markov lattice, even for the approximate direct sampling algorithm. However, if the convergence was insufficient, a sample path hit outside the Markov lattice because of the numerical error in computing transition probabilities with the floating point arithmetic. We confirmed that the generalized IPS always converged if we took small $\epsilon$ but demanded many iterations. In the following numerical experiments, we set $\epsilon=0.005$ throughout. If we could not achieve convergence by 1,000 iterations (less than 0.06%), we discarded the sample path and drew another contingency table.
Then, for $s=2,5$ and $10$, we generated a half million contingency tables using the approximate direct sampling algorithm and computed the empirical distribution $\hat{p}_{\rm D}$ of the chi-square values. For $s=1$, simple computation yields the distribution of chi-square values (19):
$p(0)=\frac{16}{37}\doteq 0.432,\quad p(8)=\frac{18}{37}\doteq 0.486,\quad p(12)=\frac{3}{37}\doteq 0.081,$ (22)
and $p(z)=0$ if $z\notin\{0,8,12\}$, and we used this.
We assessed the convergence of the Metropolis chain to the stationary distribution. For the Metropolis chain of four pairs of the burn-in and length: $(10^{3},10^{4})$, $(10^{4},10^{4})$, $(10^{5},10^{4})$ and $(10^{5},10^{5})$, respectively, we computed the empirical distribution $\hat{p}_{\rm M}$ of the chi-square values and calculated the total variation distance (20). Here, we set the expected value of each cell to $s$. We show the averages of the results obtained by repeating this procedure 10 times in Table 3. For $s=1$, we have
$\hat{p}_{M}(0)\doteq 0.427,\quad\hat{p}_{M}(8)\doteq 0.491,\quad\hat{p}_{M}(12)\doteq 0.082,\quad$
and $p(z)=0$ if $z\notin\{0,8,12\}$. The closeness to the true distribution (22) of the chi-square values demonstrates that the MLE performs a good approximation of the UMVUE. The pairs $(10^{4},10^{4})$ and $(10^{5},10^{4})$ were equally good, which implies that the Metropolis chain converged after 100,000 steps. An important observation is that the total variation distance of the pair $(10^{5},10^{5})$ decreases as $s$ increase except $s=1$ (the Monte Carlo error was small for $s=1$ since we used not the empirical distribution $\hat{p}_{D}$ but the true distribution (22)). Unlike the case of the two-way contingency tables (Table 1), the bias decrease of the MLE is evident.
Table 3: The total variation distances between the empirical distributions of chi-square values generated by the Metropolis algorithm and that by the approximate direct sampling algorithm.
(burn-in, length)
$s$ $(10^{3},10^{4})$ $(10^{4},10^{4})$ $(10^{5},10^{4})$ $(10^{5},10^{5})$
1 0.019 0.028 0.023 0.007
2 0.037 0.027 0.034 0.029
5 0.030 0.039 0.036 0.017
10 0.035 0.031 0.031 0.013
Given the above results, we drew 10,000 tables after the burn-in of 100,000 steps for the Metropolis chain. We computed the ESS of the 10,000 tables drawn by the Metropolis algorithm and drew the same number of tables using the direct sampling algorithm. Table 4 shows the averages of the results obtained by repeating this procedure 10 times. It shows the times it took to draw 10,000 tables using the Metropolis algorithm (including the burn-in), the time it took to draw the tables of the number of the ESS using the approximate direct sampling algorithm, and one table using the direct sampling algorithm. The approximate direct sampling algorithm was uniformly inefficient than the Metropolis algorithm. When the total count is 180 ($s=10$), it took about 36 times as long as the Metropolis algorithm.
Table 4: The times to draw 10,000 tables by the Metropolis algorithm, the effective sample sizes (ESS) of the 10,000 tables, the times to draw the same number of tables by the direct sampling algorithm, and the times per single table. Times are in seconds.
$s$ ESS Metropolis Direct 1 table
1 444 3.877 25.4 0.057
2 844 4.012 83.7 0.099
5 717 4.241 130.3 0.182
10 544 4.574 165.0 0.303
4 Discussion
Previous implementations of the direct sampling algorithm [14, 16] have enabled exact sampling, but depending on the model, the computational burden makes it challenging to put into practical use. In this study, we proposed the approximate algorithm applicable to any log-affine model with the expense of the bias, that asymptotically disappears. In contrast to deriving the Markov basis in the Metropolis algorithm or the connection matrix of the holonomic ideal in the exact sampling, we can implement the approximate sampling algorithm straightforwardly since it demands no prior investigations; it only needs existing methods to compute the MLE.
There is room for improvement in the approximate direct sampling algorithm proposed in this paper. In terms of accuracy, it can approach the UMVUE by reducing the bias of the MLE of the expected count, and in terms of the speed of computing the MLE, Newton’s method can be faster than the generalized iterative proportional scaling. However, these are typical tasks in statistical analysis when applying an algorithm to each model and are beyond the scope of this paper, which aims to propose an approximate sampling algorithm in its most general form.
Sampling from discrete exponential families has been a long-standing problem, but as this paper showed, we can relax it by sequentially applying the MLE instead of the UMVUE. This relaxation is a simple idea; it would not be surprising if it had existed half a century ago.
Acknowledgements
This work was supported in part by JSPS KAKENHI Grant Numbers 20K03742 and 24K06876.